OpenStudy (anonymous):

f(x)=e(x/2)-2 calculat the area bounded to f, the x-axis and the line x = 4

OpenStudy (anonymous):

Hello again.

OpenStudy (anonymous):

Ok I understand the question but I also need to know the bounds of the area.... so the top and bottom will be x-axis and the function... but what about the left and right sides?

OpenStudy (anonymous):

x=4 and x=?

OpenStudy (anonymous):

(in order right and left)

OpenStudy (anonymous):

well it is not given in the problem, but the domain = real numbers

OpenStudy (anonymous):

Well as I mentioned before f(x) has an asymptote of x=-2 so the area to the left of the y-axis is -infinity.... so this implies the left bound on x should be x=0...however I only suggested that the bound is the zero of f(x) because the first part of the question had us find it.

OpenStudy (anonymous):

Implies it should be x=0 because if the domain is all the reals then the integral of the function from -infinity to 4 = -infinity

OpenStudy (anonymous):

Im assuming this is a calculus class yes?

OpenStudy (anonymous):

When you have done problems like this before... did you ever calculate negative area? By this I mean did you ever calculate the area UNDER the x-axis? If not then the bounds should be 2 ln(2) (the zero of the function) to 4 since here the area is strictly positive. Otherwise from x=0 to x=2 ln(2) the area is finite but negative (i.e. the curve is UNDER the x-axis not above)

OpenStudy (anonymous):

If you havent already... look: http://www.wolframalpha.com/input/?i=graph+the+function+e%5E%28x%2F2%29+-+2+from+x%3D-5+to+x%3D5

OpenStudy (anonymous):

Check the other examples in your textbooks and see what they do. Do they ever have a curve BELOW the x-axis and have you calculate that area? Or is it ALWAYS area ABOVE the x-axis? (btw please note my caps are only meant to emphasize the important words)

OpenStudy (anonymous):

can i contact you if i have problems that cannot solve in the future

OpenStudy (anonymous):

sure if I am on I will help.. but what about the above questions? This a calculus class, correct? What did the other examples in your textbook do?

OpenStudy (anonymous):

yes it is. it caculate only above x-axies

OpenStudy (anonymous):

Ok brilliant

OpenStudy (anonymous):

The area under a curve is given by the integral of the function between the bounds

OpenStudy (misty1212):



OpenStudy (anonymous):

So we are looking for: \[ \int_{2 \ln(2)}^4 (e^{\frac{x}{2}} - 2 ) dx \]

OpenStudy (anonymous):

Do you know the anti-derivative of the exponential function?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

BTW hello @misty1212 your drawing is far better than any I have ever done here.... hence why I try to avoid them XD

OpenStudy (anonymous):

Ok then mayada please calculate and post the anti-derivative of: \[e^{\frac{x}{2}}-2\] and we will compare answers :D

OpenStudy (misty1212):

lol thanks!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Thanx misty1212 :) PlasmaFuzer it will be Integration by Parts right?

OpenStudy (anonymous):

No fortunately we can do this in one go... no by-parts is necessary.

OpenStudy (anonymous):

So for the function: \[f(x)=e^{ax}\]The anti-derivative is: \[F(x)=\frac{e^{ax}}{a}\]

OpenStudy (anonymous):

This website has all of a sudden become EXTREMELY laggy (facepalms) this is really frustrating

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I typed those in 10 seconds... took 2 minutes for the text to display and to post and for me to regain control of the text box.

OpenStudy (anonymous):

I have the same problem here. so iam not the only one :P

OpenStudy (anonymous):

XD yes you are not alone... my advice would be to type in a text editor and cut and paste into the box... So back to the problem... what did you calculate the anti-derivative of the function to be?